A quality-control technician is testing the concentration of muriatic acid (hydrochloric acid) to check that the concentration is within certain limits. Calculate the concentration of the hydrochloric acid if a 15.0-mL sample is titrated with standard sodium carbonate solution. The titration required 10.00 mL of 0.350 mol/L sodium carbonate to neutralize the acid.
1
Expert's answer
2019-08-12T07:40:06-0400

n(Na2CO3) = C(Na2CO3) × V(Na2CO3) = 0.350 mol/L × 0.010 L = 0.0035 mol.


0.0035 mol   X mol

    Na2CO3 + 2HCl = 2NaCl + H2O + CO2

     1 mol     2 mol


n(HCl) = X = 0.007 mol.


C(HCl) = n(HCl) / V(HCl) = 0.007 mol / 0.015 L = 0.4667 mol/L ≈ 0.467 mol/L.